The absolute temperature of a gas at which 6.79 X 10 -2 moles occupies 164 mL at 893 torr. 

1
Expert's answer
2021-07-19T04:04:38-0400

Volume of gas (V) = 164mL = 0.164 L


No of moles (n) = 6.79 × 10-2 mol = 0.0679 mol


Pressure of gas (P) = 893 torr


Gas Constant(R)= 62.36 L.Torr/mol.K


By Ideal Gas law, PV = nRT


So, T = PV/nR =


893 × 0.164 / 0.0679 × 62.36


So, the Temperature of gas(T) = 34.587 K